OpenStudy (anonymous):

How many different combinations of nine different carrots can be chosen from a bag of 20? 167,960 125,970 77,520 184,756 @iGreen

OpenStudy (anonymous):

@dan815

OpenStudy (igreen):

I think you multiply \(9 \times 8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1\).

OpenStudy (igreen):

Nvm, that's not right.

OpenStudy (igreen):

Oh, it's from a bag of 20.

OpenStudy (igreen):

I don't know. I never learned about combinations.

OpenStudy (mathmate):

combination is choosing from n different object r objects. C(n,r)=\(\large (_r^n)=\frac{n!}{(n-r)!r!}\) Here n=20, r=9.

OpenStudy (anonymous):

is this probablity